# Natural Counterfactuals With Necessary Backtracking

### v2

@@ A.4 Standard Deviation of Experimental Results
Table 8: Standard Deviation of Results on Toy $1-4$.
@@
Intersection between Ours and NB (NCO=1, NB=1) (NCO=1, NB=0) (NCO=0, NB=1) (NCO=0, NB=0)
@@
Table 10: Standard Deviation of Ablation Study on $\epsilon$
@@
Theorem $1$ from Lu et al. (2020) details identifiable counterfactuals under a certain condition:
@@
Suppose ${\bf V}_{i}$ satisfies the following structural causal model:
@@
where ${\bf U}_{i}\perp{\mathbf{PA}}_{i}$ and assume unknown $f_{i}$ is smooth and strictly monotonic w.r.t. ${\bf U}_{i}$ for fixed values of ${\mathbf{PA}}_{i}$, If we have observed ${\bf V}_{i}={\bf v}_{i}$ and ${\mathbf{PA}}_{i}=pa_{i}$, with an intervention $do({\mathbf{PA}}_{i}=pa^{*}_{i})$, the counterfactual outcome is identifiable:
@@
We use this theorem to support our proof for Theorem 4.1, which extends Theorem B.1 to a broader range of identifiable counterfactuals.
@@
We initially assume the presence of complete evidence, meaning ${\bf E}={\bf V}$. In line with Theorem B.1 from (Lu et al., 2020), counterfactual outcome of ${\bf B}_{k}\in{\bf B}$ is identifiable when given the counterfactual action ${\mathbf{PA}}_{{\bf B}_{k}}=pa_{{\bf B}_{k}}^{*}$, alongside the observable values ${\bf B}_{k}={\bf b}_{k}$ and ${\mathbf{PA}}_{{\bf B}_{k}}=pa_{{\bf B}_{k}}$, where ${\bf B}_{k}$ is any variable in ${\bf B}$ and ${\mathbf{PA}}_{{\bf B}_{k}}$ is the parent set of ${\bf B}_{k}$.
With this full evidence assumption, $pa_{{\bf B}_{k}}^{*}$ is conclusively determined by the complete evidence and the intervention $do({\bf C}={\bf c}^{*})$, whereas $pa_{{\bf B}_{k}}$ and ${\bf B}_{k}={\bf b}_{k}$ are ascertained from the full evidence. This ensures the counterfactual outcome of ${\bf B}_{k}$, a single value, is identifiable. Extending this logic, the counterfactual outcome of ${\bf B}$ as a whole, also a single value, is identifiable. Consequently, $p({\bf B}|do({\bf C}={\bf c}^{*}),{\bf E}={\bf e})$ is identifiable with a singular value as its support.
[…]
Appendix C Model Training
@@
Our study focuses on counterfactual inference and we directly use two state-of-the-art deep-learning SCM models to learn SCMs among variables using a dataset, i.e., D-SCM (Pawlowski et al., 2020) and H-SCM (Ribeiro et al., 2023). Specifically, we use code of (Ribeiro et al., 2023) containing the implementation of D-SCM and H-SCM. Take Morpho-MNIST as an example, in both two models, normalizing flows are firstly trained to learn causal mechanisms for all variables except image $x$, i.e., $(t,i)$, and a conditional VAE is used to model image $x$ given its parents $(t,i)$. For D-SCM, the conditional VAE uses normal VAE framework, while H-SCM uses hierarchical VAE structure (Maaløe et al., 2019) to better capture the distribution of images.
@@
Toy Experiments. In the case of four toy experiments, we exclusively employed normalizing flows due to the fact that all variables are one-dimensional. Our training regimen for the flow-based model spanned $2000$ epochs, utilizing a batch size of $100$ in conjunction with the AdamW optimizer (Loshchilov & Hutter, 2018). We initialized the learning rate to $1\times 10^{-3}$, set $\beta_{1}$ to $0.9$, $\beta_{2}$ to $0.9$.
@@
MorphoMNIST. We first train normalized flows to learn causal mechanisms of thickness and intensity $(t,i)$. Other hyper-parameters are similar to those of toy experiments. Then, we train two VAE-based models (D-SCM and H-SCM) to learn $x$ given $(t,i)$ respectively. The architectures of the two models are identical to (Ribeiro et al., 2023). D-SCM and H-SCM underwent training for $160$ epochs. We employed a batch size of $32$ and utilized the AdamW optimizer. The initial learning rate was set to $31e-$ 3 and underwent a linear warmup consisting of 100 steps. We set $\beta_{1}$ to $0.9$, $\beta_{2}$ to $0.9$, and applied a weight decay of $0.01$. Furthermore, we implemented gradient clipping at a threshold of $350$ and introduced a gradient update skipping mechanism, with a threshold set at $500$ based on the $L^{2}$ norm. During the testing, i.e., counterfactual inference, we test performance on both models respectively, with the normalized flows.
@@
$3$DIdentBOX. Similar to experiments on MorphoMNIST, we first train normalized flows. Compared with D-SCM, H-SCM is more powerful to model complex data like $3$DIdentBOX, of which the size of the image is $64\times 64\times 4$. Then, we train H-SCM to capture the distribution of $x$ given its parents for 500 epochs with a batch size of $32$. The hyper-parameters are the same as experiments on MorphoMNIST.
@@ Appendix D Feasible Intervention Optimization
Our loss function for Feasible Intervention Optimization is defined as:
$\mathcal{L}=\left\|an({\bf A})-an({\bf A})^{*}\right\|_{1}+w_{\epsilon}\sum_{j}[\max(\epsilon-F({\bf V}_{j}={\bf v}_{j}^{*}|pa_{j}^{*}),0)+\max(\epsilon+F({\bf V}_{j}={\bf v}_{j}^{*}|pa_{j}^{*})-1,0)]$ (7)
$*\displaystyle s.t.\quad{\bf A}={\bf a}^{*}$
Based on Theorem 4.1, Choice (3) of naturalness constraints is equivalent to Choice (2), as $F({\bf V}_{j}={\bf v}_{j}^{*}|pa_{j}^{*})=F({\bf U}_{j}={\bf u}^{*}_{j})$ with ${\bf v}_{j}^{*}=f_{j}(pa_{j}^{*},{\bf u}^{*}_{j})$. Hence, the loss can be rewritten as:
$\mathcal{L}=\left\|an({\bf A})-an({\bf A})^{*}\right\|_{1}+w_{\epsilon}\sum_{j}[\max(\epsilon-F({\bf u}_{j}^{*}),0)+\max(\epsilon+F({\bf u}_{j}^{*})-1,0)]$ (8)
$s.t.\quad{\bf u}^{*}_{\bf A}=f^{-1}_{A}({\bf a}^{*},{\mathbf{pa}}^{*}_{\bf A})$
Note that ${\bf u}_{\bf A}$ is not explicitly optimized as its value, ${\mathbf{pa}}_{\bf A}$, is determined by other noise values and the constant ${\bf a}^{*}$ through the reversible function $f^{-1}_{A}$. Consequently, ${\bf u}_{\bf A}$ is indirectly determined by other noise values and ${\bf a}^{*}$. The loss’s parameter is thus ${\bf u}_{{\mathbf{AN}}}^{*}$, which fully determines the value $an({\bf A})^{*}$ using the learned SCM, as explained in Sec. C.
In all experiments, we optimized ${\bf u}_{{\mathbf{AN}}}^{*}$ using the AdamW optimizer at a learning rate of $31e-$ 3 for $50,000$ steps. This approach’s effectiveness is validated by the MorphoMNIST experiments.
Appendix E Standard Procedure of Natural Counterfactuals
In this section, we outline our natural counterfactuals’ standard procedure, assuming full evidence. Through Feasible Intervention Optimization, we obtain the optimal value $an({\bf A})^{*}$. This identifies the changed variables, ${\bf C}$, which we target for intervention, enabling the definition of the feasible intervention $do({\bf C}={\bf c}^{*})$, with ${\bf c}^{*}$ as the post-optimization counterfactual values of ${\bf C}$. Then, we can do natural counterfactuals given the feasible intervention $do({\bf C}={\bf c}^{*})$. Accordingly, we provide a three-step standard procedure for our general natural counterfactuals:
Theorem E.1 (Natural Counterfactuals).
With the model $<{\mathcal{M}},p({\bf U})>$ and $\epsilon$, the probability $p({\bf B}_{\bf A}|{\bf e})$ of a counterfactual condition ”Given the evidence ${\bf E}={\bf e}$, if ${\bf A}$’s value is changed into ${\bf a}^{*}$ due to the feasible interventions on causally earlier variables, then ${\bf B}$,” can be predicted by the following three steps.
  • (1)
    Abduction: Given evidence ${\bf e}$, obtain updated distribution $p({\bf U}|{\bf e})$.
  • (2)
    Action: Determine feasible intervention $do({\bf C}={\bf c}^{*})$ by Feasible Intervention Optimization to achieve $change({\bf A}={\bf a}^{*})$ and modify the model into ${\mathcal{M}}_{\bf C}$ by cutting off the link between ${\bf C}$ and their parents.
  • (2)
    Prediction: Compute the probability of ${\bf B}$, the counterfactual consequence, using the updated model $<{\mathcal{M}}_{\bf C},p({\bf U}|{\bf e})>$.
@@ Appendix E Standard Procedure of Natural Counterfactuals
Appendix F Differences between Natural Counterfactuals and Non-Backtracking Counterfactuals (Pearl, 2009) or Prior-Based Backtracking Counterfactuals (von Kügelgen et al., 2022)
F.1 Differences between Non-backtracking Counterfactuals and Ours
@@
Non-backtracking counterfactuals only do a direct intervention on target variable ${\bf A}$, while our natural counterfactuals do backtracking when the direct intervention is infeasible. Notice that when the direct intervention on ${\bf A}$ is already feasible, our procedure of natural counterfactuals will be automatically distilled to the non-backtracking counterfactuals. In this sense, non-backtracking counterfactual reasoning is our special case.
F.2 Differences between Prior-Based Backtracking Counterfactuals and Ours
@@
Prior-based Backtracking Counterfactuals: These counterfactuals directly intervene on noise/exogenous variables, which can lead to unnecessary changes in the counterfactual world. Consequently, the similarity between the actual data point and its counterfactual counterpart tends to be lower. In short, prior-based backtracking counterfactuals may introduce changes that are not needed.
@@
Appendix G Observations about the Prior-Based Backtracking Counterfactuals (von Kügelgen et al., 2022)
G.1 Possibility of Gratuitous Changes

### v3

@@ A.4 Standard Deviation of Experimental Results
Table 9: Standard Deviation of Results on Toy 1 to Toy 4.
@@
Intersection between Ours and NB (NC=1, NB=1) (NC=1, NB=0) (NC=0, NB=1) (NC=0, NB=0)
@@
Table 11: Standard Deviation of Ablation Study on $\epsilon$ Using MorphoMNIST
@@
Theorem 4.1 shows that our natural counterfactuals are identifiable given certain conditions. We unitize Theorem $1$ from (Lu et al., 2020) to assistant us to prove it, shown as below.
@@
Given actual distribution $p({\mathbf{PA}}_{i},{\bf U}_{i})$, Suppose ${\bf V}_{i}$ satisfies the following function:
@@
where ${\bf U}_{i}\perp{\mathbf{PA}}_{i}$ and assume unknown $f_{i}$ is smooth and strictly monotonic w.r.t. ${\bf U}_{i}$ for fixed values of ${\mathbf{PA}}_{i}$, If we have observed ${\bf V}_{i}={\bf v}_{i}$ and ${\mathbf{PA}}_{i}=pa_{i}$, with an intervention $do({\mathbf{PA}}_{i}=pa^{*}_{i})$, where $pa^{*}_{i}$ is within the support of $p({\mathbf{PA}}_{i})$, the counterfactual outcome is identifiable:
@@
The theorem is not identical to Theorem $1$ from (Lu et al., 2020), as we incorporate additional assumptions from (Lu et al., 2020) in our proof to express it strictly. The proof is as below.
@@
Step 1. Identifiability of ${\mathbf{AN}}({\bf C})$: Given an intervention $do({\bf C}={\bf c}^{*})$ on the set of variables ${\bf C}$, where ${\mathbf{AN}}({\bf C})$ denotes the ancestors of ${\bf C}$ including ${\bf C}$ itself. By definition of the intervention and the causal graph structure, the value of ${\mathbf{AN}}({\bf C})$ under the intervention $do({\bf C}={\bf c}^{*})$ can be uniquely determined. This is because the values of ancestors and ${\bf C}$ are directly set by the intervention, making their values identifiable.
Step 2. Identifiability of ${\bf V}\setminus{\mathbf{AN}}({\bf C})$: There are two subsets in ${\bf V}\setminus{\mathbf{AN}}({\bf C})$. (1) ND: Variables in ${\bf V}\setminus{\mathbf{AN}}({\bf C})$ that are not descendants of ${\mathbf{AN}}({\bf C})$. These variables retain their values from the actual observed data since they are not influenced by the changes in ${\mathbf{AN}}({\bf C})$ due to the causal independence. (2) DS: Variables in ${\bf V}\setminus{\mathbf{AN}}({\bf C})$ that are descendants of ${\mathbf{AN}}({\bf C})$. The values of these variables might change as a result of the intervention due to their causal dependency on ${\mathbf{AN}}({\bf C})$.
Identifiability of ND: The counterfactual values of variables in ND remain as their observed values in the actual data, as they are causally independent from ${\mathbf{AN}}({\bf C})$. Thus, their values are trivially identifiable.
Identifiability of DS: The counterfactual values of variables in DS depend on the values of their parents. For variable set $\textbf{DS}_{1}$ in DS whose parents only come from $\textbf{ND}\cup{\mathbf{AN}}({\bf C})$, their counterfactual values are identifiable since, by Theorem B.1, counterfactual values of their parent variables are within the support of the joint distribution. Now, the the set of identifiable variables is $\textbf{DS}_{1}\cup\textbf{ND}\cup{\mathbf{AN}}({\bf C})$. This allows us to recursively expand the set of identifiable variables. By incorporating variables from DS whose counterfactual values are identifiable into the set $\textbf{ND}\cup{\mathbf{AN}}({\bf C})$, we iteratively enlarge this set. This process is repeated until all variables in ${\bf V}$ are included in the set of identifiable variables, thus establishing the identifiability of their values under the counterfactual scenario induced by the intervention $do({\bf C}={\bf c}^{*})$.
Conclusion: By proving the identifiability of both ${\mathbf{AN}}({\bf C})$ and all variables in ${\bf V}\setminus{\mathbf{AN}}({\bf C})$ (split into ND and DS), we establish that the entire set of endogenous variables ${\bf V}$ under the intervention $do({\bf C}={\bf c}^{*})$ is identifiable. This means the values of all variables in ${\bf V}$ can be determined uniquely from the intervention and the observed data distribution.
Appendix C Causal Model Training
@@
Our study focuses on counterfactual inference and we directly use two state-of-the-art deep-learning SCM models to learn conditional distributions among variables using a dataset, i.e., V-SCM (Pawlowski et al., 2020b) and H-SCM (Ribeiro et al., 2023). Specifically, we use code of (Ribeiro et al., 2023) containing the implementation of V-SCM and H-SCM. Take MorphoMNIST as an example, in both two models, normalizing flows are firstly trained to learn causal mechanisms for all variables except image $x$, i.e., $(t,i)$, and a conditional VAE is used to model image $x$ given its parents $(t,i)$. For V-SCM, the conditional VAE uses normal VAE framework, while H-SCM uses hierarchical VAE structure (Maaløe et al., 2019) to better capture the distribution of images.
@@
Toy Experiments. In the case of four toy experiments, we exclusively employed normalizing flows due to the fact that all variables are one-dimensional. Our training regimen for the flow-based model spanned $2000$ epochs, utilizing a batch size of $100$ in conjunction with the AdamW optimizer (Loshchilov & Hutter, 2018). We initialized the learning rate to $10^{-3}$, set $\beta_{1}$ to $0.9$, $\beta_{2}$ to $0.9$.
@@
MorphoMNIST. We first train normalized flows to learn causal mechanisms of thickness and intensity $(t,i)$. Other hyper-parameters are similar to those of toy experiments. Then, we train two VAE-based models (V-SCM and H-SCM) to learn $x$ given $(t,i)$ respectively. The architectures of the two models are identical to (Ribeiro et al., 2023). V-SCM and H-SCM underwent training for $160$ epochs. We employed a batch size of $32$ and utilized the AdamW optimizer. The initial learning rate was set to $31e-$ 3 and underwent a linear warmup consisting of 100 steps. We set $\beta_{1}$ to $0.9$, $\beta_{2}$ to $0.9$, and applied a weight decay of $0.01$. Furthermore, we implemented gradient clipping at a threshold of $350$ and introduced a gradient update skipping mechanism, with a threshold set at $500$ based on the $L^{2}$ norm. During the testing, i.e., counterfactual inference, we test performance on both models respectively, with the normalized flows.
@@
$3$DIdentBOX. Similar to experiments on MorphoMNIST, we first train normalized flows. Compared with V-SCM, H-SCM is more powerful to model complex data like $3$DIdentBOX, of which the size of the image is $64\times 64\times 4$. Then, we train H-SCM to capture the distribution of $x$ given its parents for 500 epochs with a batch size of $32$. The hyper-parameters are the same as experiments on MorphoMNIST.
All the experiments above were run on NVIDIA RTX 4090 GPUs.
@@ Appendix D Feasible Intervention Optimization
To implement a particular method, we plug the distance measure (Eqn. 4) and a naturalness constraint (we use Choice (3) in Sec. 4.1.1 in the experiments) into Eqn. 2 of the FIO framework:
$\underset{an({\bf A})^{*}}{\text{minimize}}\left\|an({\bf A})^{*}-an({\bf A})\right\|_{1}$ (8)
$s.t.\quad{\bf A}={\bf a}^{*},$
According to the properties of reversible functions in the causal model, endogenous value $an({\bf A})^{*}$ is reversible to noise values ${\bf u}^{*}_{{\mathbf{AN}}({\bf A})}$ of ${\mathbf{AN}}({\bf A})^{*}$’s corresponding noise variables. Hence, optimizing endogenous value is equivalent to optimizing noise value. Then we have:
$\underset{{\bf u}_{{\mathbf{AN}}({\bf A})}^{*}}{\text{minimize}}\sum_{{\bf u}_{j}^{*}\in{\bf u}_{{\mathbf{AN}}({\bf A})}^{*}}|\hat{f}(pa^{*}_{\bf A},{\bf u}_{j}^{*})-{\bf v}_{j}|$ (9)
$s.t.\quad{\bf u}^{*}_{\bf A}=\hat{f}^{-1}_{\bf A}(pa^{*}_{\bf A},{\bf a}^{*}),$
$\quad\epsilon<F({\bf u}_{j}^{*})<1-\epsilon,\forall{\bf U}_{j}\in{\bf U}_{{\mathbf{AN}}({\bf A})}.$
We finally use the Lagrangian method (Boyd & Vandenberghe, 2004) optimize our objective loss to get the counterfactual value given actual value and expected change ${\bf A}={\bf a}^{*}$ as below:
$\mathcal{L}({\bf u}^{*}_{{\mathbf{AN}}({\bf A})})=\sum_{{\bf u}_{j}^{*}\in{\bf u}_{{\mathbf{AN}}({\bf A})}^{*}}|\hat{f}_{j}({\bf u}_{j}^{*},pa_{j}^{*})-{\bf v}_{j}|+w_{\epsilon}\sum_{{\bf u}_{j}^{*}\in{\bf u}_{{\mathbf{AN}}({\bf A})}^{*}}[\max(\epsilon-F({\bf u}_{j}^{*}),0)+\max(\epsilon+F({\bf u}_{j}^{*})-1,0)]$ (10)
$s.t.\quad{\bf u}^{*}_{\bf A}=\hat{f}^{-1}_{\bf A}(pa^{*}_{\bf A},{\bf a}^{*})$
where the optimization parameters are the counterfactual noise values of ${\bf A}$’s ancestors, ${\bf u}^{*}_{{\mathbf{AN}}({\bf A})}$, and the function $\max(\cdot)$ returns the maximum value between two given values. The first term represents the measure of distance between two distinct worlds, while the second term enforces the constraint of $\epsilon$-natural generation. Here, the constant hyperparameter $w_{\epsilon}$ serves to penalize noise values situated in the tails of noise distributions. For simplicity, we use ${\bf A}$ as subscript as indicator of terms related to ${\bf A}$, instead of number subscript. Notice that, in order to ensure hard constraint ${\bf A}={\bf a}^{*}$, ${\bf A}$’s noise value ${\bf u}^{*}_{\bf A}$ is not optimized explicitly, since the value ${\mathbf{pa}}^{*}_{\bf A}$ is fully determined by ${\bf a}^{*}$ and other noise values.
Hyper-parameters for Optimization and Judgment for Non-solution Cases. The loss’s parameter is thus ${\bf u}_{{\mathbf{AN}}}^{*}$, which fully determines the value $an({\bf A})^{*}$ using the pretrained causal model, as explained in Sec. C. In all experiments, we optimized ${\bf u}_{{\mathbf{AN}}}^{*}$ using the AdamW optimizer at a learning rate of $10^{-3}$ for $50,000$ steps. This approach’s effectiveness is validated by the MorphoMNIST experiments, as shown in Table 5. As $\hat{f}$ is not perfectly invertible in practice, it is important to ensure that $|pa^{*}_{\bf A}-\hat{f}_{\bf A}({\bf u}^{*}_{\bf A},{\bf a}^{*})|$ (where ${\bf u}^{*}_{\bf A}=\hat{f}^{-1}_{\bf A}(pa^{*}_{\bf A},{\bf a}^{*})$) remains sufficiently small, ideally on the order of $10^{-8}$ according to our experience, to satisfy the constraint condition in Eqn. 10. In our experience, this deviation $|pa^{*}_{\bf A}-\hat{f}_{\bf A}({\bf u}^{*}_{\bf A},{\bf a}^{*})|$ is typically sufficiently small throughout most of the optimization process. If not, please adjust hyperparameters, such as the learning rate and $w_{\epsilon}$, to achieve the desired accuracy. Assuming $|pa^{*}_{\bf A}-\hat{f}_{\bf A}({\bf u}^{*}_{\bf A},{\bf a}^{*})|$ is sufficiently small after optimization, if $\sum_{{\bf u}_{j}^{*}\in{\bf u}_{{\mathbf{AN}}({\bf A})}^{*}}[\max(\epsilon-F({\bf u}_{j}^{*}),0)+\max(\epsilon+F({\bf u}_{j}^{*})-1,0)]>0$ holds, this indicates that the case does not have a solution.
Computational Complexity of FIO. The complexity scales linearly with both the size of the causal graph, specifically the number of ancestors of $A$, and the dataset size. Formally, for one counterfactual query, the overall complexity is $O(KPT)$, where $K$ is the number of ancestors of $A$, $P$ is the number of parameters in the neural networks, and $T$ is the number of optimization steps. If we use each instance of a whole dataset as evidence and answer one counterfactual query for each instance, the overall complexity becomes $O(KPTM)$, where $M$ is the number of data points in the dataset. This linear scaling indicates that FIO is reasonably scalable, making it suitable for large causal graphs and datasets.
@@
Appendix E Differences between Natural Counterfactuals and Non-Backtracking Counterfactuals (Pearl, 2009) or Prior-Based Backtracking Counterfactuals (von Kügelgen et al., 2022)
E.1 Differences between Non-backtracking Counterfactuals and Ours
@@
Non-backtracking counterfactuals only do a direct intervention on ${\bf A}$, while our natural counterfactuals do backtracking when the direct intervention is infeasible. Notice that when the direct intervention on ${\bf A}$ is already feasible, our procedure of natural counterfactuals will be automatically distilled to the non-backtracking counterfactuals. In this sense, non-backtracking counterfactual reasoning is our special case.
E.2 Differences between Prior-Based Backtracking Counterfactuals and Ours
@@
Prior-based Backtracking Counterfactuals: These counterfactuals directly intervene on noise/exogenous variables, which can lead to unnecessary changes in the counterfactual world. [^5] To avoid misleading readers, we clarify why we apply the concept of intervention to exogenous variables. Prior-based backtracking counterfactuals do not appeal to interventions in the usual Pearlian sense, for those are restricted to endogenous variables and are associated with breaking endogenous mechanisms. However, formally speaking, changing the values of exogenous variables is analogous to intervening on exogenous variables, in the sense that they change those variables without affecting the mechanisms for other variables in the model (the invariance of the mechanisms for other variables is for many the crucial hallmark of an intervention). In fact, some authors explicitly advocate applying the notion of intervention to exogenous variables as well (Vandenburgh, 2023). Consequently, the similarity between the actual data point and its counterfactual counterpart tends to be lower. In short, prior-based backtracking counterfactuals may introduce changes that are not needed.
@@
Appendix F Observations about the Prior-Based Backtracking Counterfactuals (von Kügelgen et al., 2022)
F.1 Possibility of Gratuitous Changes